Question 1

(a) The rates of growth of money supply is 10%, of velocity of money circulation 1%, of real GDP 3%, what is the inflation rate?

(b) The nominal interest rate is 7%, the inflation rate is 5%, what is the real interest rate?
Answer #1

a.

Velocity of money forumula

V=(P*Y)/M

V= velocity of money

M=money supply growth

P=price level or inflation

Y=real GDP

Since it is given that The rates of growth of money supply is 10%, of velocity of money circulation 1%, of real GDP 3%,

the inflation rate (P)=(V*M)/Y

P=(1*10)/3

P=10/3

P=3.33%

Hence the inflation rate will be 3.33%

b.

The nominal interest rate is 7%, the inflation rate is 5%,

Real interest rate=Nominal interest rate - inflation rate

=7%-5%

=2%